Solve using substitution:
3x+5y =-6 ; x=5y+3
I have solved that Y= 3/4 but im not sure what to do with the fraction to figure x out. Thank You. Angela

Answer by jim_thompson5910(35256)

You should have gotten y = -3/4

Once you have y, you can use it to find x


x=5y+3


x=5(-3/4)+3


x=-15/4+3


x=-15/4+12/4


x=(-15+12)/4


x=-3/4


So the solutions are x = -3/4 and y = -3/4